# On the Estimation of bivariate Conditional Transition Rates

### v1

@@ 5 Bivariate Conditional Nelson-Aalen and Aalen-Johansen Estimation
on $C_{m}$. Referring to the fact that $\boldsymbol{Q}_{z,\boldsymbol{ij}}^{c}(\tau)$ is bounded and that $\boldsymbol{N}_{z,\boldsymbol{ij}}^{(n)}(\boldsymbol{\tau})\xrightarrow[n\rightarrow\infty]{a.s.}\boldsymbol{Q}_{z,\boldsymbol{ij}}^{c}(\boldsymbol{\tau})$, $\mathbb{P}(C_{m})=1$ for $m$ large enough. All in all, we get the almost surely uniform convergence of $\boldsymbol{N}_{z,\boldsymbol{ij}}^{(n)}(\boldsymbol{t})$.
@@
Similar to the approach with $\boldsymbol{P}_{z,\boldsymbol{i}}(\boldsymbol{t})$ and $\boldsymbol{I}_{z,\boldsymbol{i}}^{(n)}(\boldsymbol{t})$, we can now separate $C_{z,\boldsymbol{i}}(\boldsymbol{t})$ and $\boldsymbol{C}_{z,\boldsymbol{i}}^{(n)}(\boldsymbol{t})$ into four summands that are monotonously increasing and bounded on a set with probability $1$. Thus, with the same two-dimensional Glivenko-Cantelli arguments we used before and the strong law of large numbers, we get the almost sure uniform convergence of $\boldsymbol{C}_{z,\boldsymbol{i}}^{(n)}(\boldsymbol{t})$ to $\boldsymbol{C}_{z,\boldsymbol{i}}(\boldsymbol{t})$. With equations (5.5), (5.4), the uniform convergence of the univariate estimators and the law of large numbers, we get the almost sure uniform convergence of $\boldsymbol{I}_{z,\boldsymbol{i}}^{(n)}$ to $\boldsymbol{P}_{z,\boldsymbol{i}}^{c}$ on $(s,\boldsymbol{\tau}]$. ∎
@@
for $\boldsymbol{t}\in(s,\boldsymbol{\tau}]$. Thus, $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}^{(\epsilon)}(\boldsymbol{t})\xrightarrow[\epsilon\rightarrow 0]{a.s.}\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}(\boldsymbol{t}),$ for all $\boldsymbol{t}\in(s,\boldsymbol{\tau}]$. Now $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}^{(\epsilon)}$ and $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}$ are both monotonously increasing and bounded on $(s,\boldsymbol{\tau}]$. With arguments similar to the two-dimensional proof of the Gliveno-Cantelli theorem the pointwise convergence can be extended to
@@ 6 Uniform Convergence in Actuarial Estimation
Let $\kappa$ be a càdlàg function that describes the value development of a savings account. We assume that $\kappa$ is strictly positive so that the corresponding discounting function $1/\kappa$ exists. The random variable
$Y^{+}:=\;\int\displaylimits_{(s,T]}\frac{\kappa(s)}{\kappa(u)}B(\mathrm{d}u)$ (6.1)
describes the discounted accumulated future cash flow seen from time $s$. In order to ensure integrability, we generally assume that $1/\kappa$ is bounded on $[0,T]$.
Classical Markov modelling uses payment functions that allow payments to depend on the current state of the insured and on current jumps from one state to another. A classical cash flow that includes both of these option is the so-called one-dimensional canonical representation.
@@
A stochastic process $A$ is said to have a one-dimensional canonical representation if there exist real-valued functions $(A_{i})_{i}$ on $[0,\infty)$ with finite variation on finite intervals and measurable functions $(a_{ij})_{ij:i\neq j}$, which also have finite variation on finite intervals such that
$A(t)=\sum_{i\in\mathcal{Z}}\;\int\displaylimits_{(s,t]}I_{i}(u^{-})A_{i}(\mathrm{d}u)+\sum_{i,j\in\mathcal{Z}\atop i\neq j}\;\int\displaylimits_{(s,t]}a_{ij}(u^{-})N_{ij}(\mathrm{d}u),\quad t\geq s.$ (6.2)
@@
A stochastic process $A$ is said to have a two-dimensional canonical representation if there exist real-valued functions $(A_{i})_{i}$ on $[0,\infty)$ which are a difference of two non-decreasing upper continuous functions, real-valued functions $(A_{ij})_{ij}$ on $[0,\infty)^{2}$ which are a difference of two non-decreasing upper-continuous functions, and measurable and bounded real-valued functions $(a_{ikl})_{ikl}$, $(a_{ijkl})_{ijkl}$ on $[0,\infty)^{2}$ which are also the difference of two non-decreasing upper-continuous functions, such that
$\begin{split}\boldsymbol{A}(\boldsymbol{t})&=\sum_{i,j\in\mathcal{Z}}\;\;\int\displaylimits_{(s,\boldsymbol{t}]}I_{i}(u_{1}^{-})I_{j}(u_{2}^{-})A_{ij}(\mathrm{d}u_{1},\mathrm{d}u_{2})\\ &\quad+\sum_{\begin{subarray}{c}i,k,l\in\mathcal{Z}\\ k\neq l\end{subarray}}\;\int\displaylimits_{(s,\boldsymbol{t}]}I_{i}(u_{1}^{-})a_{ikl}(u_{1}^{-},u_{2}^{-})A_{i}(\mathrm{d}u_{1})N_{kl}(\mathrm{d}u_{2})\\ &\quad+\sum_{\begin{subarray}{c}i,j,k,l\in\mathcal{Z}\\ i\neq j,k\neq l\end{subarray}}\;\int\displaylimits_{(s,\boldsymbol{t}]}a_{ijkl}(u_{1}^{-},u_{2}^{-})N_{ij}(\mathrm{d}u_{1})N_{kl}(\mathrm{d}u_{2}),\quad\boldsymbol{t}\geq s.\end{split}$ (6.3)
@@
$\boldsymbol{V}_{z}:=\mathbb{E}\Big{[}\int\displaylimits_{(s,T]^{2}}\boldsymbol{A}(\mathrm{d}\boldsymbol{u})\Big{|}\xi=z\Big{]},$
$V_{z}:=\mathbb{E}\Big{[}\int\displaylimits_{(s,T]^{2}}A(\mathrm{d}u)\Big{|}\xi=z\Big{]}.$
Bathke & Christiansen [4] showed
$\begin{split}\boldsymbol{V}_{z}&=\sum_{i,j\in\mathcal{Z}}\;\int\displaylimits_{(s,T]^{2}}\boldsymbol{P}_{z,ij}(\boldsymbol{u}^{-})A_{ij}(\mathrm{d}\boldsymbol{u})\\ &\quad+\sum_{\begin{subarray}{c}i,k,l\in\mathcal{Z}\\ k\neq l\end{subarray}}I_{i}(s)\int\displaylimits_{(s,T]^{2}}a_{ikl}(u_{1}^{-},u_{2}^{-})A_{i}(\mathrm{d}u_{1}){P}_{z,k}(u_{2}^{-})\Lambda_{z,kl}(\mathrm{d}u_{2})\\ &\quad+\sum_{\boldsymbol{i}\in\mathcal{Z},\boldsymbol{j}\in\mathcal{Z}_{\neq}}\;\;\int\displaylimits_{(s,T]}\;\int\displaylimits_{(s,T]\times(s,u_{1})}{a}_{j_{2}i_{1}j_{1}}(u_{1}^{-},u_{2}^{-})\boldsymbol{P}_{z,\boldsymbol{i}}(u_{2}^{-},u_{3}^{-})\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}(\mathrm{d}u_{2},\mathrm{d}u_{3})A_{j_{2}}(\mathrm{d}u_{1})\\ &\quad+\sum_{\boldsymbol{i},\boldsymbol{j}\in\mathcal{Z}_{\neq}}\;\int\displaylimits_{(s,T]^{2}}a_{i_{1}j_{1}i_{2}j_{2}}(u_{1}^{-},u_{2}^{-})\boldsymbol{P}_{z,\boldsymbol{i}}(\boldsymbol{u}^{-})\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}(\mathrm{d}\boldsymbol{u}^{-}).\end{split}$ (6.4)
@@
$V_{z}$ $=\sum_{i\in\mathcal{Z}}\int\displaylimits_{(s,T]}P_{z,i}(u^{-})A_{i}(\mathrm{d}u)+\sum_{\begin{subarray}{c}i,j\in\mathcal{Z}\\ i\neq j\end{subarray}}\int\displaylimits_{(s,T]}a_{ij}(u^{-})P_{z,i}(u^{-})\Lambda_{z,ij}(\mathrm{d}u).$
These $\boldsymbol{V}_{z}$ and $V_{z}$ can be understood as a functionals $\boldsymbol{V}_{z}({P}_{z},\boldsymbol{P}_{z},\Lambda_{z},\boldsymbol{\Lambda}_{z})$ and $V_{z}({P}_{z},\Lambda_{z})$ of the uni- and bivariate conditional transition rates and transition probabilities. For the purposes of reserving in the insurer’s balance sheet, we now need to estimate these conditional expected values. In practice, we first estimate the conditional transition rates and probabilities as seen in the previous section and then plug them into the functionals $\boldsymbol{V}_{z}$ and $V_{z}$. This works because of the following theorems.
@@
For a stochastic process that has a one-dimensional canonical cash-flow representation, the functional $V_{z}$ is continuous in sup-norm in all $2$ arguments, meaning that the convergence in sup-norm in the domain results in convergence in the codomain.
$\lim_{\epsilon\rightarrow 0}\lim_{n\rightarrow\infty}\Big{|}V_{z}({P}_{z}^{(n,\epsilon)},\Lambda_{z}^{(n,\epsilon)})-V_{z}({P}_{z},\Lambda_{z})\Big{|}\stackrel{{\scriptstyle a.s.}}{{=}}0.$
The proof of this theorem will follow at the end of this section.
@@
For a stochastic process that has a two-dimensional canonical cash-flow representation, the functional $\boldsymbol{V}_{z}$ is continuous in sup-norm in all $4$ arguments, meaning that the convergence in sup-norm in the domain results in convergence in the codomain.
@@
$\lim_{\epsilon\rightarrow 0}\lim_{n\rightarrow\infty}\Big{|}\boldsymbol{V}_{z}({P}_{z}^{(n,\epsilon)},\boldsymbol{P}_{z}^{(n,\epsilon)},\Lambda_{z}^{(n,\epsilon)},\boldsymbol{\Lambda}_{z}^{(n,\epsilon)})-\boldsymbol{V}_{z}({P}_{z},\boldsymbol{P}_{z},\Lambda_{z},\boldsymbol{\Lambda}_{z})\Big{|}\stackrel{{\scriptstyle a.s.}}{{=}}0.$
The proof of this theorem will follow at the end of this section.
@@
    Plug these estimators in the functionals $\boldsymbol{V}_{z}$ and $V_{z}$.
@@
We will show the continuity of $V_{z}$ with the help of Lemma A.3 and especially the fact that $B(F,G)\mapsto\int F(s)\mathrm{d}G(s)$ is Hadamard differentiable. For this differentiability to apply, we need bounded variation of all the integrators and integrands in equation (6.4). For the first integral, we have bounded variation in $n$ of $\boldsymbol{P}_{z}^{(n,\epsilon)}$ and $\boldsymbol{P}_{z}$ as we have seen in the proof of Theorem 5.7 with the use of equation (2.3), additionally we have that $A_{ij}$ is of bounded variation because it was defined as the difference of two non-decreasing upper continuous functions.
For the second summand, we need $\int\displaylimits_{(s,T]}a_{ikl}(u_{1}^{-},u_{2}^{-})A_{i}(\mathrm{d}u_{1})$ to be of bounded variation in $u_{2}$. For that we use the fact, that $a_{ikl}$ and $A_{i}$ are both functions that can be written as the difference of two non-decreasing functions $a^{+}_{ikl}$, $a^{-}_{ikl}$, $A_{i}^{+}$ and $A_{i}^{-}$. Plugging all of these in, we can bound the variation by
$\Big{\|}\int\displaylimits_{(s,T]}a_{ikl}(u_{1}^{-},u_{2}^{-})A_{i}(\mathrm{d}u_{1})\Big{\|}_{v_{1}}^{(1)}\leq$ $\int\displaylimits_{(s,T]}a_{ikl}^{+}(u_{1}^{-},T)-a_{ikl}^{+}(u_{1}^{-},s)A_{i}^{+}(\mathrm{d}u_{1})$
$\;+\int\displaylimits_{(s,T]}a_{ikl}^{-}(u_{1}^{-},T)-a_{ikl}^{-}(u_{1}^{-},s)A_{i}^{+}(\mathrm{d}u_{1})$
$\;+\int\displaylimits_{(s,T]}a_{ikl}^{+}(u_{1}^{-},T)-a_{ikl}^{+}(u_{1}^{-},s)A_{i}^{-}\mathrm{d}u_{1})$
$\;+\int\displaylimits_{(s,T]}a_{ikl}^{-}(u_{1}^{-},T)-a_{ikl}^{-}(u_{1}^{-},s)A_{i}^{-}(\mathrm{d}u_{1}).$
Additionally, $P_{z}^{(n,\epsilon)}$ and $P_{z}$ are of uniform bounded variation because of equation 2.3. For the last two summands, we use the same ideas and the fact, that $A_{j}$, $a_{ijk}$, and $a_{ijkl}$ are all the difference of two non-decreasing functions. Thus, we proofed the continuity of $V_{z}$ and with that found a way to estimate the conditional expected value of a stochastic process with a two-dimensional canonical cash flow representation. ∎

### v2

@@ 5 Bivariate Conditional Nelson-Aalen and Aalen-Johansen Estimation
Remark 5.12.
We introduce the idea that one would follow to prove asymptotic normality, i.e. weak convergence to a tight Gaussian process of the introduced estimators. To prove the asymptotic normality of $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}^{(n,\epsilon)}(\boldsymbol{t})$ and $\boldsymbol{P}_{z}^{(n,\epsilon)}(\boldsymbol{t})$ one uses the Hadamard differentiability of the equation (5.1) interpreted as a functional and ideas introduced in the proof of Theorem 4.5 in Bladt and Furrer [7], which are easily generalised to two dimensions, for the demonstration of the asymptotic normality of $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}^{(n,\epsilon)}(\boldsymbol{t})$. Both of these ideas are based on the asymptotic normality of $\boldsymbol{N}^{(n)}_{z,\boldsymbol{ij}}(\boldsymbol{t})$ and $\boldsymbol{I}_{z,\boldsymbol{i}}^{(n)}(\boldsymbol{t})$, where the second also bases on the first with the help of equation (2.3). The asymptotic normality of $\boldsymbol{N}^{(n)}_{z,\boldsymbol{ij}}(\boldsymbol{t})$ follows with the bracketing central limit theorem, Theorem 2.11.9 in [30]. The convergence via the bracketing central limit theorem relies on the bracketing number, which is the minimum number of sets $P_{i}$ in the partition of $[s,\tau]^{2}$ such that
$\mathbb{E}[\sup_{\boldsymbol{t},\boldsymbol{s}\in P_{i}}|N_{j_{1}k_{1}}(t_{1}\wedge R)N_{j_{2}k_{2}}(t_{2}\wedge R)-N_{j_{1}k_{1}}(s_{1}\wedge R)N_{j_{2}k_{2}}(s_{1}\wedge R)|^{2}]<\mu^{2}$
for $\mu$ close to zero. This can be reduced to a problem of one-dimensional partitions, by applying the equality $ab-dc=a(b-c)+c(a-d)$, which allows us to work with differences in one dimension rather than two. This reduces the problem of the bracketing number significantly, since the jump processes are all monotonically increasing and bounded in expectation. From here, one uses the classical arguments for one-dimensional jump processes to follow the asymptotic normality. This conveys the main idea needed for the asymptotic normality of conditional bivariate transition rates and transition probabilities. Since this paper is more concerned with the statistical application of these estimators, a detailed proof is beyond the scope of this paper.
@@
on $C_{m}$. Referring to the fact that $\boldsymbol{Q}_{z,\boldsymbol{ij}}^{c}(\tau)$ is bounded and that $\boldsymbol{N}_{z,\boldsymbol{ij}}^{(n)}(\boldsymbol{\tau})\xrightarrow[n\rightarrow\infty]{a.s.}\boldsymbol{Q}_{z,\boldsymbol{ij}}^{c}(\boldsymbol{\tau})$, $\mathbb{P}(C_{m})=1$ for $m$ large enough. All in all, we get the almost sure uniform convergence of $\boldsymbol{N}_{z,\boldsymbol{ij}}^{(n)}(\boldsymbol{t})$.
@@
Similar to the approach with $\boldsymbol{P}_{z,\boldsymbol{i}}(\boldsymbol{t})$ and $\boldsymbol{I}_{z,\boldsymbol{i}}^{(n)}(\boldsymbol{t})$, we can now separate $C_{z,\boldsymbol{i}}(\boldsymbol{t})$ and $\boldsymbol{C}_{z,\boldsymbol{i}}^{(n)}(\boldsymbol{t})$ into four summands that are monotonically increasing and bounded on a set with probability $1$. Thus, with the same two-dimensional Glivenko-Cantelli arguments we used before and the strong law of large numbers, we get the almost sure uniform convergence of $\boldsymbol{C}_{z,\boldsymbol{i}}^{(n)}(\boldsymbol{t})$ to $\boldsymbol{C}_{z,\boldsymbol{i}}(\boldsymbol{t})$. With equations (5.5), (5.4), the uniform convergence of the univariate estimators and the law of large numbers, we get the almost sure uniform convergence of $\boldsymbol{I}_{z,\boldsymbol{i}}^{(n)}$ to $\boldsymbol{P}_{z,\boldsymbol{i}}^{c}$ on $(s,\boldsymbol{\tau}]$. ∎
@@
for $\boldsymbol{t}\in(s,\boldsymbol{\tau}]$. Thus, $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}^{(\epsilon)}(\boldsymbol{t})\xrightarrow[\epsilon\rightarrow 0]{a.s.}\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}(\boldsymbol{t}),$ for all $\boldsymbol{t}\in(s,\boldsymbol{\tau}]$. Now $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}^{(\epsilon)}$ and $\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}$ are both monotonically increasing and bounded on $(s,\boldsymbol{\tau}]$. With arguments similar to the two-dimensional proof of the Gliveno-Cantelli theorem the pointwise convergence can be extended to
@@ 6 Uniform Convergence in Actuarial Estimation
Similar to the general multi-state framework, we define $Z=(Z(t))_{t\geq s},$ as the state process of the insured with the state indicator process $(I_{i})_{i\in\mathcal{Z}}$ and the transition counting process $(N_{ij})_{i,j\in\mathcal{Z}}$. The landmark $\xi=Z(s)$, similar to the information model used in Markov modelling in the actuarial context, is realistic in application, but different conditioning is also feasible.
@@
Several models have been proposed to estimate the expected value of these future cash flows. The Markov model only allows for restrictive payment functions in the future cash flow $B(t)$ and requires the Markov assumption. The more general semi-Markov model allows for the estimation of cash flows with options as policyholder behaviour, see Ahmad et al. in [3]. The disadvantage of this model is that it still introduces systematic model risk because we cannot verify the required Markov assumption and that the estimation is computationally expensive. Therefore, modelling without the Markov assumption in a non-Markov model is the more general way to model future cash flows. This has been done for scaled cash flows by Christiansen and Furrer in [11]. There they use an approach that uses contract-dependent estimates to account for policyholder behaviour. The disadvantage is that, in practice, re-estimating for different insurance contracts is time-consuming and introduces other problems, particularly in terms of data usage and storage.
The conditional bivariate transition and transition probabilities introduced in this paper, allow for an estimation procedure that is independent of the design of the insurance contract and allows for a wide variety of payment functions. In practice, this variability allows a service provider to estimate these rates and probabilities once with its combined data and then distribute them to the insurance companies without distributing the data. This is already good practice, for example, with the German Actuarial Association and its semi-Markov tables.
We now introduce the different cash flows $B(t)$ that we use to model the cash flow of insurance contracts.
Classical Markov modelling uses payment functions that allow payments to depend on the current state of the insured and on current jumps from one state to another. A classical cash flow that includes both of these option is a cash flow with a so-called one-dimensional canonical representation.
@@
A stochastic process $B$ is said to have a one-dimensional canonical representation if there exist real-valued functions $(B_{i})_{i}$ on $[0,\infty)$ with finite variation on finite intervals and measurable functions $(b_{ij})_{ij:i\neq j}$, which also have finite variation on finite intervals such that
$B(t)=\sum_{i\in\mathcal{Z}}\;\int\displaylimits_{(s,t]}I_{i}(u^{-})B_{i}(\mathrm{d}u)+\sum_{i,j\in\mathcal{Z}\atop i\neq j}\;\int\displaylimits_{(s,t]}b_{ij}(u^{-})N_{ij}(\mathrm{d}u),\quad t\geq s.$ (6.1)
@@
A stochastic process $B$ is said to have a two-dimensional canonical representation if there exist real-valued functions $(B_{i})_{i}$ on $[0,\infty)$ which are a difference of two non-decreasing upper continuous functions, real-valued functions $(B_{ij})_{ij}$ on $[0,\infty)^{2}$ which are a difference of two non-decreasing upper-continuous functions, and measurable and bounded real-valued functions $(b_{ikl})_{ikl}$, $(b_{ijkl})_{ijkl}$ on $[0,\infty)^{2}$ which are also the difference of two non-decreasing upper-continuous functions, such that
$\begin{split}\boldsymbol{B}(\boldsymbol{t})&=\sum_{i,j\in\mathcal{Z}}\;\;\int\displaylimits_{(s,\boldsymbol{t}]}I_{i}(u_{1}^{-})I_{j}(u_{2}^{-})B_{ij}(\mathrm{d}u_{1},\mathrm{d}u_{2})\\ &\quad+\sum_{\begin{subarray}{c}i,k,l\in\mathcal{Z}\\ k\neq l\end{subarray}}\;\int\displaylimits_{(s,\boldsymbol{t}]}I_{i}(u_{1}^{-})b_{ikl}(u_{1}^{-},u_{2}^{-})B_{i}(\mathrm{d}u_{1})N_{kl}(\mathrm{d}u_{2})\\ &\quad+\sum_{\begin{subarray}{c}i,j,k,l\in\mathcal{Z}\\ i\neq j,k\neq l\end{subarray}}\;\int\displaylimits_{(s,\boldsymbol{t}]}b_{ijkl}(u_{1}^{-},u_{2}^{-})N_{ij}(\mathrm{d}u_{1})N_{kl}(\mathrm{d}u_{2}),\quad\boldsymbol{t}\geq s.\end{split}$ (6.2)
@@
$\boldsymbol{A}_{z}(t):=\mathbb{E}\Big{[}\int\displaylimits_{(s,t]^{2}}\boldsymbol{B}(\mathrm{d}\boldsymbol{u})\Big{|}\xi=z\Big{]},$
$A_{z}(t):=\mathbb{E}\Big{[}\int\displaylimits_{(s,t]}B(\mathrm{d}u)\Big{|}\xi=z\Big{]}.$
In practice, a positive probability landmark should be used to condition the cash flows. These cash flows with a two-dimensional canonical representation have already been studied by Bathke & Christiansen in [5]. They showed
$\begin{split}\boldsymbol{A}_{z}(t)&=\sum_{i,j\in\mathcal{Z}}\;\int\displaylimits_{(s,t]^{2}}\boldsymbol{P}_{z,ij}(\boldsymbol{u}^{-})B_{ij}(\mathrm{d}\boldsymbol{u})\\ &\quad+\sum_{\begin{subarray}{c}i,k,l\in\mathcal{Z}\\ k\neq l\end{subarray}}I_{i}(s)\int\displaylimits_{(s,t]^{2}}b_{ikl}(u_{1}^{-},u_{2}^{-})B_{i}(\mathrm{d}u_{1}){P}_{z,k}(u_{2}^{-})\Lambda_{z,kl}(\mathrm{d}u_{2})\\ &\quad+\sum_{\boldsymbol{i}\in\mathcal{Z},\boldsymbol{j}\in\mathcal{Z}_{\neq}}\;\;\int\displaylimits_{(s,t]}\;\int\displaylimits_{(s,t]\times(s,u_{1})}{b}_{j_{2}i_{1}j_{1}}(u_{1}^{-},u_{2}^{-})\boldsymbol{P}_{z,\boldsymbol{i}}(u_{2}^{-},u_{3}^{-})\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}(\mathrm{d}u_{2},\mathrm{d}u_{3})B_{j_{2}}(\mathrm{d}u_{1})\\ &\quad+\sum_{\boldsymbol{i},\boldsymbol{j}\in\mathcal{Z}_{\neq}}\;\int\displaylimits_{(s,t]^{2}}b_{i_{1}j_{1}i_{2}j_{2}}(u_{1}^{-},u_{2}^{-})\boldsymbol{P}_{z,\boldsymbol{i}}(\boldsymbol{u}^{-})\boldsymbol{\Lambda}_{z,\boldsymbol{ij}}(\mathrm{d}\boldsymbol{u}^{-}).\end{split}$ (6.3)
@@
$A_{z}(t)$ $=\sum_{i\in\mathcal{Z}}\int\displaylimits_{(s,t]}P_{z,i}(u^{-})B_{i}(\mathrm{d}u)+\sum_{\begin{subarray}{c}(i,j)\in\mathcal{Z}_{\neq}\end{subarray}}\int\displaylimits_{(s,t]}b_{ij}(u^{-})P_{z,i}(u^{-})\Lambda_{z,ij}(\mathrm{d}u).$
These $\boldsymbol{A}_{z}$ and $A_{z}$ can be understood as a functionals $\boldsymbol{A}_{z}({P}_{z},\boldsymbol{P}_{z},\Lambda_{z},\boldsymbol{\Lambda}_{z})$ and $A_{z}({P}_{z},\Lambda_{z})$ of the uni- and bivariate conditional transition rates and transition probabilities. For the purposes of reserving in the insurer’s balance sheet, we now need to estimate these conditional expected values. In practice, we first estimate the conditional transition rates and probabilities as seen in the previous section and then plug them into the functionals $\boldsymbol{A}_{z}$ and $A_{z}$. This works because of the following theorems.
@@
For a stochastic process that has a one-dimensional canonical cash flow representation, the functional $A_{z}$ is continuous in sup-norm in all arguments, meaning that the convergence in sup-norm in the domain results in convergence in the codomain. Additionally,
$\lim_{\epsilon\rightarrow 0}\lim_{n\rightarrow\infty}\Big{|}A_{z}({P}_{z}^{(n,\epsilon)},\Lambda_{z}^{(n,\epsilon)})(t)-A_{z}({P}_{z},\Lambda_{z})(t)\Big{|}\stackrel{{\scriptstyle a.s.}}{{=}}0,\quad\forall t>s.$
The proof of this theorem follows at the end of this section.
@@
For a stochastic process that has a two-dimensional canonical cash flow representation, the functional $\boldsymbol{A}_{z}$ is continuous in sup-norm in all $4$ arguments, meaning that the convergence in sup-norm in the domain results in convergence in the codomain.
@@
$\lim_{\epsilon\rightarrow 0}\lim_{n\rightarrow\infty}\Big{|}\boldsymbol{A}_{z}({P}_{z}^{(n,\epsilon)},\boldsymbol{P}_{z}^{(n,\epsilon)},\Lambda_{z}^{(n,\epsilon)},\boldsymbol{\Lambda}_{z}^{(n,\epsilon)})(t)-\boldsymbol{A}_{z}({P}_{z},\boldsymbol{P}_{z},\Lambda_{z},\boldsymbol{\Lambda}_{z})(t)\Big{|}\stackrel{{\scriptstyle a.s.}}{{=}}0,\quad\forall t>s.$
The proof of this theorem follows at the end of this section.
@@
    Plug these estimators in the functionals $\boldsymbol{A}_{z}$ and $A_{z}$.
@@
We show the theorem for $t=T$, for all $t<T$ the proof works the same. We start by showing the continuity of $\boldsymbol{A}_{z}$ with the help of Lemma A.3 and especially the fact that $B(F,G)\mapsto\int F(s)\mathrm{d}G(s)$ is Hadamard differentiable. For this differentiability to apply, we need bounded variation of all the integrators and integrands in equation (6.3). For the first integral, we have bounded variation in $n$ of $\boldsymbol{P}_{z}^{(n,\epsilon)}$ and $\boldsymbol{P}_{z}$ as we have seen in the proof of Theorem 5.9 with the use of equation (2.3), additionally we have that $B_{ij}$ is of bounded variation because it was defined as the difference of two non-decreasing upper continuous functions.
For the second summand, we need $\int\displaylimits_{(s,T]}b_{ikl}(u_{1}^{-},u_{2}^{-})B_{i}(\mathrm{d}u_{1})$ to be of bounded variation in $u_{2}$. For that we use the fact, that $b_{ikl}$ and $B_{i}$ are both functions that can be written as the difference of two non-decreasing functions $b^{+}_{ikl}$, $b^{-}_{ikl}$, $B_{i}^{+}$ and $B_{i}^{-}$. Plugging all of these in, we can bound the variation by
$\Big{\|}\int\displaylimits_{(s,T]}b_{ikl}(u_{1}^{-},u_{2}^{-})B_{i}(\mathrm{d}u_{1})\Big{\|}_{v_{1}}^{(1)}\leq$ $\int\displaylimits_{(s,T]}b_{ikl}^{+}(u_{1}^{-},T)-b_{ikl}^{+}(u_{1}^{-},s)B_{i}^{+}(\mathrm{d}u_{1})$
$\;+\int\displaylimits_{(s,T]}b_{ikl}^{-}(u_{1}^{-},T)-b_{ikl}^{-}(u_{1}^{-},s)B_{i}^{+}(\mathrm{d}u_{1})$
$\;+\int\displaylimits_{(s,T]}b_{ikl}^{+}(u_{1}^{-},T)-b_{ikl}^{+}(u_{1}^{-},s)B_{i}^{-}\mathrm{d}u_{1})$
$\;+\int\displaylimits_{(s,T]}b_{ikl}^{-}(u_{1}^{-},T)-b_{ikl}^{-}(u_{1}^{-},s)B_{i}^{-}(\mathrm{d}u_{1}).$
Additionally, $P_{z}^{(n,\epsilon)}$ and $P_{z}$ are of uniform bounded variation, which can be seen similarly to $\boldsymbol{P}_{z}^{(n,\epsilon)}$ and $\boldsymbol{P}_{z}$ with equation 2.3. For the last two summands, we use the same ideas and the fact, that $B_{j}$, $b_{ijk}$, and $b_{ijkl}$ are all the difference of two non-decreasing functions. Thus, we have proved the continuity of $A_{z}$ and with that found a way to estimate the conditional expected value of a stochastic process with a two-dimensional canonical cash flow representation. ∎